OpenStudy (cedric):

Factor each trinomial below. Please show your work and check your answer. (1 point each) 1.x2 – 8x + 15 2.a2 – a – 20 3.a2 – 5a – 20 4.a2 + 12ab + 27b2 5.2a2 + 30a + 100

OpenStudy (anonymous):

same procedure for all you look for 2 numbers(example 1) whoseproduct is 15 and whose sum is -8=>-3,-5 x^2-5x-3x+15=x(x-5)-3(x-5)=(x-5)(x-3)

OpenStudy (cedric):

oh ok thanks man!

OpenStudy (cedric):

imma try 2

OpenStudy (anonymous):

ok

OpenStudy (cedric):

5 and 4 make 20, but idk how to make it equal a variable lol

OpenStudy (anonymous):

you want -20 not 20

OpenStudy (cedric):

ohhh -5 and 4

OpenStudy (anonymous):

good

OpenStudy (cedric):

so thats all i have to right? is -5 and 4 ?

OpenStudy (anonymous):

then complete factorization as i did

OpenStudy (cedric):

idk how to do that

OpenStudy (jiteshmeghwal9):

hint to do factorization: break the middle term such that their sum becomes the middle term & product becomes the side terms

OpenStudy (jiteshmeghwal9):

now look one example

OpenStudy (jiteshmeghwal9):

\[\Huge{\color{red}{x^2-8x+15}}\]\[\Huge{\color{green}{\implies{x^2-5x-3x+15}}}\]\[\Huge{\color{green}{\implies x(x-5)-3(x-5)}}\]\[\Huge{\color{green}{\implies (x-3)(x-5)}}\]

OpenStudy (jiteshmeghwal9):

gt it @cedric ??? Do other example similarly like that :)

OpenStudy (cedric):

ok lemme try

OpenStudy (jiteshmeghwal9):

sure! take ur time

OpenStudy (anonymous):

i think that the 3rd one can't be factorized,are you sure its -5a and not -8a

OpenStudy (cedric):

so for the x's i use 5- and 4?

OpenStudy (jiteshmeghwal9):

it depends on on the equation

OpenStudy (cedric):

the second one

OpenStudy (jiteshmeghwal9):

ohh ! yes u r right @cedric :)

OpenStudy (jiteshmeghwal9):

-5a+4a

OpenStudy (jiteshmeghwal9):

s wht do u gt @cedric ??

OpenStudy (jiteshmeghwal9):

so*

OpenStudy (jiteshmeghwal9):

@cedric ???????????????????????????????????????????????????????????????????